# Characterisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v2

@@ 3. Some properties of automorphisms and their actions on $G$ and ${\mathrm{Sub}}^{p}_{G}$
The notation is inspired by the fact that the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$ defined by Dani and McCrudden in [13]. The latter consists of connected Lie groups which admit a (real) linear representation with discrete kernel. Proposition 2.5 of [13] shows that $G$ is in class $\mathcal{C}$ if and only if for the radical $R$ of $G$, the intersection of $\overline{[R,R]}$ with the center of $G$ has no non-trivial compact subgroup. Therefore class ${\mathcal{C}^{\prime}}$ is a subset of class $\mathcal{C}$. Note that any connected Lie group $G$ admits a maximal compact connected central subgroup $M$, and $G/M$ has no non-trivial compact central subgroup and hence it belongs to class ${\mathcal{C}^{\prime}}$. We may recall a result of Dani in [12] that if $G$ belongs to class ${\mathcal{C}^{\prime}}$, then ${\rm Aut}(G)$ is almost algebraic, i.e. it is an algebraic subgroup of ${\rm GL}({\mathfrak{g}})$, and in particular it has finitely many connected components.
It is known that $T\in{\rm Aut}(G)$ acts distally on $G$ if and only if ${\rm d\,}T$ acts distally on the Lie algebra ${\mathfrak{g}}$ of $G$, and the latter statement is equivalent to the statement that all the eigenvalues of ${\rm d\,}T$ have absolute value 1 (cf [2]). It is also known that if $G$ admits an automorphism $T$ such that ${\rm d\,}T$ has no eigenvalue of absolute value 1 (equivalently, if $T$ is expansive; see [38], and also [7]), then $G$ must be nilpotent (this is well-known, see e.g. [38]). We will need the following lemma which is slightly stronger. The lemma may be known, but we give a short proof for the sake of completeness.
@@
Note that $C(T)$ is a connected nilpotent subgroup of $G$ and it is simply connected if it is closed in $G$ (see [43], see also [21]). Both $C(T)$ and $C(T^{-1})$ are trivial if and only if $T$ acts distally on $G$ (cf.  [35], Theorem 4.1). Note also that $C(T)$ is trivial if and only if all the eigenvalues of ${\rm d\,}T$ have absolute value greater than or equal to 1 ([35, 2]).
@@
    The contraction groups $C(T)$ and $C(T^{-1})$ are closed, simply connected and nilpotent.
@@
We note the following useful proposition which will enable us to work on the quotient Lie group which belongs to class ${\mathcal{C}^{\prime}}$.
@@
Let $T\in{\rm Aut}(G)$ and let $\pi:G\to G/K$ be the natural projection. Note that $\overline{T}(\pi(H))=\pi(T(H))$ for any $H\in{\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G/K}$ is $\overline{T}$ invariant. Let ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=H\}$ which is a compact $T$-invariant subset of ${\rm Sub}^{p}_{G}$ and homeomorphic to ${\rm Sub}^{p}_{G/K}$ (cf. Lemma 2.4). Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then the $T$-action on ${\mathcal{S}}(K)$ is distal. Let $L_{1},L_{2}\in{\rm Sub}^{p}_{G/K}$ be such that for some $\{n_{k}\}\subset{\mathbb{Z}}$, ${\overline{T}}^{n_{k}}(L_{i})\to L$, for some $L\in{\rm Sub}^{p}_{G/K}$, $i=1,2$. Then there exist $H_{1},H_{2}\in{\mathcal{S}}(K)$, such that $\pi(H_{i})=L_{i}$ and $T^{n_{k}}(H_{i})\to\pi^{-1}(L)$ in ${\mathcal{S}}(K)$, $i=1,2$ (cf. Lemma 2.4 (4)). As $T$ acts distally on ${\mathcal{S}}(K)$, we get that $H_{1}=H_{2}$, and hence that $L_{1}=L_{2}$. Thus $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/K}$. ∎
@@
In a simply connected nilpotent group, all one-parameter subgroups are closed and there is a one-to-one correspondence between non-trivial (closed) one-parameter subgroups of $G$ and one-dimensional subspaces of ${\mathfrak{g}}$. The correspondence is via the exponential map $\exp:{\mathfrak{g}}\to G$ with $\log$ as its inverse. Namely, for $\{x_{t}\}$, $x_{t}=\exp tX$, for some $X\in{\mathfrak{g}}$, and $\log x_{t}=tX$, $t\in{\mathbb{R}}$. Suppose $\{H_{n}\}\subset{\mathcal{P}}_{1}(G)$ is such that $H_{n}\to H$ in ${\rm Sub}_{G}$. Then $\log H_{n}\in{\mathcal{P}}_{1}({\mathfrak{g}})$ and using Lemma 2.1 and the continuity of the exponential map and its inverse, it is easy to show that $\{\log H_{n}\}$ converges to (say) $B$ in ${\rm Sub}^{p}_{\mathfrak{g}}$. Since ${\mathfrak{g}}$ is isomorphic to ${\mathbb{R}}^{n}$ for some $n$, ${\rm Sub}^{p}_{\mathfrak{g}}={\mathcal{P}}_{1}({\mathfrak{g}})$ which is closed and we have that $B\in{\mathcal{P}}_{1}({\mathfrak{g}})$. Then $H=\exp B\in{\mathcal{P}}_{1}(G)$, and we get that ${\mathcal{P}}_{1}(G)$ is closed in ${\rm Sub}_{G}$, i.e. ${\mathcal{P}}_{1}(G)={\rm Sub}^{p}_{G}$. Therefore, $H_{n}\to H$ in ${\rm Sub}^{p}_{G}$ implies that $\log H_{n}\to\log H$ in ${\rm Sub}^{p}_{\mathfrak{g}}$.
Conversely, if $L_{n}\to L$ in ${\mathcal{P}}_{1}({\mathfrak{g}})$, using Lemma 2.1 and the fact that $\exp$ is a homeomorphism from ${\mathfrak{g}}$ to $G$, it is easy to see that $\exp L_{n}\to\exp L$ in ${\mathcal{P}}_{1}(G)$. Now the exponential map defines a homeomorphism from ${\mathcal{P}}_{1}({\mathfrak{g}})$ to ${\mathcal{P}}_{1}(G)$ and $T$ (resp. ${\rm d\,}T$) defines a homeomorphism of ${\mathcal{P}}_{1}(G)$ (resp. ${\mathcal{P}}_{1}({\mathfrak{g}})$) such that $\exp\circ\,{\rm d\,}T=T\circ\exp$. Therefore, it is easy to see that $T$ acts distally on ${\mathcal{P}}_{1}(G)={\rm Sub}^{p}_{G}$ if and only if ${\rm d\,}T$ acts distally on ${\mathcal{P}}_{1}({\mathfrak{g}})={\rm Sub}^{p}_{\mathfrak{g}}$. ∎
@@ 4. Proofs of the main results
Conversely, for $G={\mathbb{R}}^{d}$ and $T\in{\mathcal{K}}\mathcal{D}$, with a compact set ${\mathcal{K}}\subset{\rm GL}(d,{\mathbb{R}})$ and the center ${\mathcal{D}}$ of ${\rm GL}(d,{\mathbb{R}})$, the action of $T$ on ${\rm Sub}^{p}_{G}$ is same as that of $S$ on ${\rm Sub}^{p}_{G}$ for some $S\in{\mathcal{K}}$ since $\mathcal{D}$ acts trivially on ${\rm Sub}^{p}_{G}$. As the compact subgroups of ${\rm Aut}(G)$ act distally on any closed invariant subspace of ${\rm Sub}_{G}$ (${\rm Sub}^{p}_{G}$ in this case), we have that $T$ acts distally on ${\rm Sub}^{p}_{G}$. This completes the proof of Theorem 1.1 (1).
@@
As $G$ is simply connected and nilpotent, by Lemma 3.6, ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$. Then from (1) above, we have that ${\rm d\,}T\in{\mathcal{K}}\{r{\rm Id}\mid r\neq 0\}$ for some compact subgroup ${\mathcal{K}}$ of ${\rm GL}({\mathfrak{g}})$ and ${\rm Id}$ is the identity map on ${\mathfrak{g}}$. Here, ${\rm d\,}T=\psi\,\circ\,r\,{\rm Id}$ for some $r\in{\mathbb{R}}\setminus\{0\}$ and $\psi\in{\mathcal{K}}$. If $r=\pm 1$, then ${\rm d\,}T^{2}\in{\mathcal{K}}$ and $T$ generates a relatively compact subgroup of ${\rm Aut}(G)$. If possible suppose $r\neq\pm 1$. Since ${\rm Aut}(G)$ is almost algebraic, there exists $n\in{\mathbb{N}}$ such that ${\rm d\,}T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathbb{R}}\times\mathbb{T}^{k}$ of ${\rm Aut}(G)^{0}$ (which is identified with its image in ${\rm GL}({\mathfrak{g}})$), where $\mathbb{T}^{k}=(\mathbb{S}^{1})^{k}$ for some $k\in{\mathbb{N}}$. Therefore, $\psi^{n}\,\circ\,r^{n}\,{\rm Id}\in{\mathbb{R}}\times\mathbb{T}^{k}\subset{\rm Aut}(G)^{0}$. Hence both $\psi^{n}$ and $r^{n}\,{\rm Id}\in{\rm Aut}(G)^{0}$. But $r^{n}\,{\rm Id}$ is not a Lie algebra automorphism of ${\mathfrak{g}}$ as $r^{n}\neq 1$; otherwise $r^{n}[X,Y]=r^{n}{\rm Id}[X,Y]=[r^{n}{\rm Id}(X),r^{n}{\rm Id}(Y)]=r^{2n}[X,Y]$, then $[X,Y]=0$ for all $X,Y\in{\mathfrak{g}}$, which leads to a contradiction as $G$ is not abelian. Thus, $r=\pm 1$ and $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse statement is obvious. Thus (2) holds if $G$ is nilpotent. ∎
@@
Let $G$ be a connected Lie group which is not a vector group. Let $M$ be the largest compact connected central subgroup of $G$. Suppose $T\in{\rm Aut}(G)$ be such that it acts distally on $M$. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then we want to show that $T$ acts distally on $G$.
@@
Step 2: Now suppose $G$ is not a simply connected nilpotent group. By Theorem 4.1 of [35], it is enough to prove that the contraction groups $C(T)$ and $C(T^{-1})$ are trivial. As $T$ acts distally on the maximal compact connected central subgroup $M$, by Lemma 3.4, both $C(T)$ and $C(T^{-1})$ are closed, simply connected and nilpotent. Suppose $C(T)$ is non-trivial. If possible, suppose $C(T)$ is non-abelian. As it is $T$-invariant and $T|_{C(T)}$ acts distally on ${\rm Sub}^{p}_{C(T)}$, we get from Step 1 that $T|_{C(T)}$ acts distally on $C(T)$, a contradiction. Now suppose $C(T)$ is abelian. Then $C(T)$ is isomorphic to ${\mathbb{R}}^{d}$ as $C(T)$ is simply connected. By Theorem 1.1 (1), $T_{1}=T|_{C(T)}=\phi\circ\alpha\,I$ for some $\phi$ contained in a compact subgroup of ${\rm GL}(d,{\mathbb{R}})$ and $\alpha\in{\mathbb{R}}$, $10<|\alpha|<$ 1; here $I$ denotes the identity matrix in ${\rm GL}(d,{\mathbb{R}})$. We may replace $T$ by $T^{2}$ (and $T_{1}$ by $T_{1}^{2}$), and assume that $\alpha>0$. Here, $G\neq C(T)$ as $G$ is not a vector group.
@@
Suppose $G$ is nilpotent but not simply connected. Then $M$ as above is non-trivial. Since $T$ acts distally on $M$, by Lemma 2.5 of [2], we have that $T^{k}|_{M}$ is unipotent for some $k\in{\mathbb{N}}$, i.e. ${\rm d\,}T$ has an eigenvalue which is a root of unity. If $G$ is not nilpotent, by Lemma 3.2, ${\rm d\,}T$ has an eigenvalue which is a root of unity. We may replace $T$ by a suitable power and assume that 1 is an eigenvalue of ${\rm d\,}T$.
Let $X\in{\mathfrak{g}}$, $X\neq 0$ be such that ${\rm d\,}T(X)=X$ and let $A=\{\exp tX\}$, a one-parameter subgroup, it is either closed or its closure $\bar{A}$ is compact. Since $T|_{A}={\rm Id}$, we have that $T|_{\bar{A}}={\rm Id}$ and it follows that $\bar{A}$ normalises $C(T)$ and $\bar{A}\cap C(T)=\{e\}$. Let $H=\overline{A\ltimes C(T)}$. Then $H$ is solvable. As $A$ normalises $C(T)$, so does $H$. Now $H/C(T)$ is connected and abelian as it is the closure of the image of the one-parameter subgroup $A$ in it. Then $H/C(T)$ is either isomorphic to ${\mathbb{R}}$ or it is compact. In the first case it follows that $H=A\ltimes C(T)$. In the second case, since $C(T)$ is a vector group, $H=K\ltimes C(T)$ for some (maximal) compact subgroup $K\subset H$ (cf. [25], Lemma 3.7). If $A$ (resp. $\bar{A}$) is compact, then $H=A\ltimes C(T)$ (resp. $H=\bar{A}\ltimes C(T)$). Here, $H$ is a closed $T$-invariant subgroup of $G$ and $T(h)\in h\,C(T)$ for all $h\in H$.
Step 3: Suppose $H/C(T)$ is isomorphic to ${\mathbb{R}}$. Then so is $A$ and $H=A\ltimes C(T)$. Let $U,V$ be open neighbourhoods of the identity $e$ in $H$ such that $V^{2}\subset U$ and every element of $U$ is contained in a one-parameter subgroup. Let $x=ac$ for some $a\in A\cap V$ and $c\in C(T)\cap V$ with $a\neq e$ and $c\neq e$. Then $x\in V^{2}\subset U$ and $T(x)=xy$ for $y=c^{-1}T(c)\in C(T)$ and $y\neq e$ as $c\in C(T)$ and $c\neq e$. Now $x$ is contained in a one-parameter subgroup (say) $\{x_{t}\}$ in $H$ as $x=x_{1}$. Then $T(x_{t})=x_{t}y_{t}$, $y_{t}\in C(T)$, $t\in{\mathbb{R}}$ and $y_{1}=y\neq e$. Here, $\{x_{t}\}$ is closed as $H$ is simply connected and solvable. As ${\rm Sub}^{p}_{H}$ is compact, we get that $T^{n_{k}}(\{x_{t}\})\to B$ for some $B\in{\rm Sub}^{p}_{H}$ and some unbounded sequence $\{n_{k}\}\subset{\mathbb{N}}$. By Lemma 2.2, $B\neq\{e\}$. Let $b\in B\setminus\{e\}$. Then $T^{n_{k}}(x_{t_{k}})\to b$ for some sequence $\{t_{k}\}\subset{\mathbb{R}}$. Here, $\{x_{t}\}C(T)/C(T)=H/C(T)$ is isomorphic to ${\mathbb{R}}$. Since $x_{t_{k}}C(T)\to b\,C(T)$ in $H/C(T)$, we get that $\{t_{k}\}$ is bounded. Passing to a subsequence if necessary, we may assume that $t_{k}\to t_{0}$. Then $x_{t_{k}}\to x_{t_{0}}$ and $y_{t_{k}}\to y_{t_{0}}$. As all the eigenvalues of $T|_{C(T)}=T_{1}=\phi\circ\alpha\,I$ have absolute value $\alpha$ with $10<\alpha<$ 1, we get that $I-T_{1}$ is invertible in ${\rm GL}(d,{\mathbb{R}})$, where $I$ is the identity matrix in ${\rm GL}(d,{\mathbb{R}})$. Using the addition as the group operation in ${\mathbb{R}}^{d}$, and the fact that $T_{1}^{n}(z)\to 0$ for all $z\in{\mathbb{R}}^{d}$, we get that
$T^{n_{k}}(x_{t_{k}})=x_{t_{k}}\left(\sum_{i=0}^{n_{k-1}}T_{1}^{i}(y_{t_{k}})\right)\to x_{t_{0}}(I-T_{1})^{-1}(y_{t_{0}})=b$
(as $n_{k}\to\infty$), where $T_{1}^{0}=I$, the identity in ${\rm GL}(d,{\mathbb{R}})$. Note that
$I+T_{1}(I-T_{1})^{-1}=(I-T_{1})(I-T_{1})^{-1}+T_{1}(I-T_{1})^{-1}=(I-T_{1})^{-1}.$
Therefore,
$T(b)=x_{t_{0}}y_{t_{0}}T_{1}(I-T_{1})^{-1}(y_{t_{0}})=x_{t_{0}}(I+T_{1}(I-T_{1})^{-1})(y_{t_{0}})=b.$
Thus $T(B)=B$. Now we get that $T^{n_{k}}(T(\{x_{t}\}))\to T(B)=B$ and hence that $T(\{x_{t}\})=\{x_{t}\}$ since $T$ acts distally on ${\rm Sub}^{p}_{G}$. As $T$ is an automorphism, we have that $T(x_{t})=x_{st}$ for some $s\in{\mathbb{R}}\setminus\{0\}$, and $xy=x_{1}y_{1}=x_{s}$. Therefore, $x_{s-1}=y_{1}=y\neq e$ and $s\neq 1$. Then $x_{s-1}\in C(T)$, and $\{x_{t}\}C(T)/C(T)$ is compact. This leads to a contradiction as $H=\{x_{t}\}C(T)$ and $H/C(T)$ is isomorphic to ${\mathbb{R}}$.
Step 4: Now we have that $H/C(T)$ is compact and $H=K\ltimes C(T)$, where $K$ is a maximal compact subgroup of $H$. We show that $T(x)=x$ for all $x\in K$. We know that $T(x)\in x\,C(T)$ for all $x\in K$. Since $K$ is connected and abelian, the set $F$ of roots of unity in $K$ is dense in $K$ and it is enough to show that $T(x)=x$ for all $x\in F$. If possible, suppose there exists $x\in F$ such that $T(x)\neq x$. Let $m\in{\mathbb{N}}$ be the order of $x$, i.e. $x^{m}=e$, and $m\neq 1$. Let $\{x_{t}\}$ be a one-parameter subgroup in $K$ such that $x_{1}=x$. As $x_{1}^{m}=e$, $\{x_{t}\}$ is closed and isomorphic to ${\mathbb{S}}^{1}$ and $T(x_{t})=x_{t}y_{t}$, where $y_{t}\in C(T)$, $t\in{\mathbb{R}}$ and $y_{1}=x^{-1}T(x)\neq e$. There exists an unbounded sequence $\{n_{k}\}\subset{\mathbb{N}}$ such that $T^{n_{k}}(\{x_{t}\})\to B$ for some $B\in{\rm Sub}^{p}_{H}$. Note that $B\neq\{e\}$ (cf. Lemma 2.2). Let $b\in B$ be such that $b\neq e$. Then $T^{n_{k}}(x_{t_{k}})\to b$ for some $\{t_{k}\}\subset{\mathbb{R}}$. Here, $\{x_{t}\}$ is periodic i.e. $x_{mn}=e$ and $x_{m+t}=x_{t}$ for all $n\in{\mathbb{Z}}$ and $t\in{\mathbb{R}}$, where $m$ is the order of $x_{1}=x$. Hence we can choose $\{t_{k}\}$ such that it is bounded. Then $\{x_{t_{k}}\}$, and hence, $\{y_{t_{k}}\}$ is bounded. Now arguing as in Step 3, we get that $T(b)=b$. Therefore, $T(B)=B$, and hence $T^{n_{k}}(T(\{x_{t}\}))\to B$. Since $T$ acts distally on ${\rm Sub}^{p}_{G}$, we get that $T(\{x_{t}\})=\{x_{t}\}$. As $T$ is an automorphism and $\{x_{t}\}$ is isomorphic to ${\mathbb{S}}^{1}$, we get that $T(x_{t})=x_{t}$ or $T(x_{t})=x_{t}^{-1}=x_{-t}$. In either case, we get that $y_{1}\in K\cap C(T)=\{e\}$. This leads to a contradiction. Therefore, $T(x)=x$, for all $x\in F$, and hence for all $x\in K$.
Step 5: We now show that $K$ centralises $C(T)$ and $H=K\times C(T)$. We have that $T|_{K}={\rm Id}$. Let $K_{c}=c^{-1}Kc$, $c\in C(T)$. Then $K_{c}$ is a compact connected abelian group and $H=K_{c}\ltimes C(T)$ and $T(x)\in x\,C(T)$ for all $x\in K_{c}$. Therefore, arguing as in Step 4 for $K_{c}$ instead of $K$, we get that $T|_{K_{c}}={\rm Id}$, i.e. $c^{-1}xc=T(c^{-1}xc)=T(c^{-1})x\,T(c)$ for all $c\in C(T)$ and $x\in K$. Thus $c\,T(c^{-1})=(I-T_{1})(c)$ centralises $K$, $c\in C(T)$, where $T_{1}=T|_{C(T)}$. As $I-T_{1}$ is invertible in $GL(d,{\mathbb{R}})$, we get that $C(T)$ centralises $K$. Thus $H=K\times C(T)$.

### v3

@@ 3. Some properties of automorphisms and their actions on $G$ and ${\mathrm{Sub}}^{p}_{G}$
The notation is inspired by the fact that the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$ defined by Dani and McCrudden in [13]. The class $\mathcal{C}$ consists of connected Lie groups which admit a (real) linear representation with discrete kernel. Proposition 2.5 of [13] shows that $G$ is in class $\mathcal{C}$ if and only if for the radical $R$ of $G$, the intersection of $\overline{[R,R]}$ with the center of $G$ has no non-trivial compact subgroup. Therefore, the class ${\mathcal{C}^{\prime}}$ is a subset of the class $\mathcal{C}$. Note that any connected Lie group $G$ admits a maximal compact connected central subgroup $M$, and $G/M$ has no non-trivial compact connected central subgroup and hence it belongs to class ${\mathcal{C}^{\prime}}$. We may recall a result of Dani in [12] that if $G$ belongs to class ${\mathcal{C}^{\prime}}$, then ${\rm Aut}(G)$ is almost algebraic, i.e. it is a(n open) subgroup of finite index in an algebraic subgroup of ${\rm GL}({\mathfrak{g}})$, and in particular, it has finitely many connected components.
It is known that $T\in{\rm Aut}(G)$ acts distally on $G$ if and only if ${\rm d\,}T$ acts distally on the Lie algebra ${\mathfrak{g}}$ of $G$, and the latter statement is equivalent to the statement that all the eigenvalues of ${\rm d\,}T$ have absolute value 1 (cf. [2]). It is also known that if $G$ admits an automorphism $T$ such that ${\rm d\,}T$ has no eigenvalue of absolute value 1 (equivalently, if $T$ is expansive; see [38], and also [7]), then $G$ must be nilpotent (this is well-known, see e.g. [38]). We will need the following lemma which is slightly stronger. The lemma may be known, but we will give a short proof for the sake of completeness.
@@
Note that $C(T)$ is a connected nilpotent subgroup of $G,$ and if it is closed in $G$, then it is simply connected (see [43], see also [21]). Both $C(T)$ and $C(T^{-1})$ are trivial if and only if $T$ acts distally on $G$ (cf.  [35], Theorem 4.1). Note also that $C(T)$ is trivial if and only if all the eigenvalues of ${\rm d\,}T$ have absolute value greater than or equal to 1 (cf. [35, 2]).
@@
    The contraction groups $C(T)$ and $C(T^{-1})$ are closed, simply connected and nilpotent subgroups of $G$.
@@
We note the following useful proposition which will enable us to work on a quotient Lie group which belongs to class ${\mathcal{C}^{\prime}}$.
@@
Let $T\in{\rm Aut}(G)$ and let $\pi:G\to G/K$ be the natural projection. Here, $\overline{T}(\pi(H))=\pi(T(H))$ for any $H\in{\rm Sub}_{G}$, and ${\rm Sub}^{p}_{G/K}$ is $\overline{T}$ invariant. Let ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=H\}$. It is a compact $T$-invariant subset of ${\rm Sub}^{p}_{G}$ and it is homeomorphic to ${\rm Sub}^{p}_{G/K}$ under the map induced by $\pi$ on ${\mathcal{S}}(K)$ (cf. Lemma 2.4 (4)). Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then the $T$-action on ${\mathcal{S}}(K)$ is distal. As $\pi\circ T=\overline{T}\circ\pi$, we get that $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/K}$. ∎
@@
In a simply connected nilpotent Lie group, all one-parameter subgroups are closed and there is a one-to-one correspondence between non-trivial (closed) one-parameter subgroups of $G$ and one-dimensional subspaces of ${\mathfrak{g}}$. The correspondence is via the exponential map $\exp:{\mathfrak{g}}\to G$ with $\log$ as its inverse. Namely; for $\{x_{t}\}$, $x_{t}=\exp tX$, for some $X\in{\mathfrak{g}}$, and $\log x_{t}=tX$, $t\in{\mathbb{R}}$. Also the exponential map induces a homeomorphism from ${\mathcal{P}}_{1}({\mathfrak{g}})\to{\mathcal{P}}_{1}(G)$. As $\exp\circ\,{\rm d\,}T=T\circ\exp$, it follows that $T$ acts distally on ${\mathcal{P}}_{1}(G)={\rm Sub}^{p}_{G}$ if and only if ${\rm d\,}T$ acts distally on ${\mathcal{P}}_{1}({\mathfrak{g}})={\rm Sub}^{p}_{\mathfrak{g}}$. ∎
@@ 4. Proofs of the main results
Conversely, for $G={\mathbb{R}}^{d}$ and $T\in{\mathcal{K}}{\mathcal{D}}$, with a compact subgroup ${\mathcal{K}}$ of ${\rm GL}(d,{\mathbb{R}})$ and the center ${\mathcal{D}}$ of ${\rm GL}(d,{\mathbb{R}})$, the action of $T$ on ${\rm Sub}^{p}_{G}$ is same as that of $S$ on ${\rm Sub}^{p}_{G}$ for some $S\in{\mathcal{K}}$ since ${\mathcal{D}}$ acts trivially on ${\rm Sub}^{p}_{G}$. As the compact subgroups of ${\rm Aut}(G)$ act distally on any closed invariant subspace of ${\rm Sub}_{G}$ (${\rm Sub}^{p}_{G}$ in this case), we have that $T$ acts distally on ${\rm Sub}^{p}_{G}$. This completes the proof of Theorem 1.1 (1).
@@
As $G$ is simply connected and nilpotent, by Lemma 3.6, ${\rm d\,}T$ acts distally on ${\rm Sub}^{p}_{\mathfrak{g}}$. Then from (1) above, we have that ${\rm d\,}T\in{\mathcal{K}}\{r\,{\rm Id}\mid r\neq 0\}$ for some compact subgroup ${\mathcal{K}}$ of ${\rm GL}({\mathfrak{g}})$ and ${\rm Id}$ is the identity map on ${\mathfrak{g}}$. Here, ${\rm d\,}T=\psi\,\circ\,r\,{\rm Id}$ for some $r\in{\mathbb{R}}\setminus\{0\}$ and $\psi\in{\mathcal{K}}$. If $r=\pm 1$, then ${\rm d\,}T^{2}\in{\mathcal{K}}$ and $T$ generates a relatively compact subgroup of ${\rm Aut}(G)$. If possible suppose $r\neq\pm 1$. Since ${\rm Aut}(G)$ is almost algebraic, there exists $n\in{\mathbb{N}}$ such that ${\rm d\,}T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathbb{R}}\times\mathbb{T}^{k}$ of ${\rm Aut}(G)^{0}$ (which is identified with its image in ${\rm GL}({\mathfrak{g}})$), where $\mathbb{T}^{k}=(\mathbb{S}^{1})^{k}$ for some $k\in{\mathbb{N}}$. Therefore, $\psi^{n}\,\circ\,r^{n}\,{\rm Id}\in{\mathbb{R}}\times\mathbb{T}^{k}\subset{\rm Aut}(G)^{0}$. Hence both $\psi^{n}$ and $r^{n}\,{\rm Id}\in{\rm Aut}(G)^{0}$. But $r^{n}\,{\rm Id}$ is not a Lie algebra automorphism of ${\mathfrak{g}}$ as $r^{n}\neq 1$; otherwise $r^{n}[X,Y]=r^{n}{\rm Id}[X,Y]=[r^{n}{\rm Id}(X),r^{n}{\rm Id}(Y)]=r^{2n}[X,Y]$, and it would imply that $[X,Y]=0$ for all $X,Y\in{\mathfrak{g}}$, which leads to a contradiction as $G$ is not abelian. Thus, $r=\pm 1$ and $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. The converse statement is obvious. Thus (2) holds if $G$ is nilpotent. ∎
@@
Let $G$ be a connected Lie group which is not a vector group. Let $M$ be the largest compact connected central subgroup of $G$. Suppose $T\in{\rm Aut}(G)$ be such that it acts distally on $M$. Suppose $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then we want to show that $T$ acts distally on $G$. The assertion holds trivially if $G=M$.
@@
Step 2: Now suppose $G$ is not a simply connected nilpotent group. By Theorem 4.1 of [35], it is enough to prove that the contraction groups $C(T)$ and $C(T^{-1})$ are trivial. As $T$ acts distally on the maximal compact connected central subgroup $M$, by Lemma 3.4, both $C(T)$ and $C(T^{-1})$ are closed, simply connected and nilpotent subgroups. Suppose $C(T)$ is non-trivial. If possible, suppose $C(T)$ is non-abelian. As it is $T$-invariant and $T|_{C(T)}$ acts distally on ${\rm Sub}^{p}_{C(T)}$, we get from Step 1 that $T|_{C(T)}$ acts distally on $C(T)$, a contradiction. Thus $C(T)$ is abelian and $C(T)$ is isomorphic to ${\mathbb{R}}^{d}$ as it is simply connected. By Theorem 1.1 (1), $T_{1}=T|_{C(T)}=\phi\circ\alpha\,I$ for some $\phi$ contained in a compact subgroup of ${\rm GL}(d,{\mathbb{R}})$ and $\alpha\in{\mathbb{R}}$, $10<|\alpha|<$ 1; here $I$ denotes the identity matrix in ${\rm GL}(d,{\mathbb{R}})$, replacing $T$ by $T^{2}$, we may assume that $\alpha>0$.
@@
Step 3: Suppose $G$ is nilpotent but not simply connected. Then $M$ as above is non-trivial. Since $T$ acts distally on $M$, by Lemma 2.5 of [2], we have that $T^{k}|_{M}$ is unipotent for some $k\in{\mathbb{N}}$, i.e. ${\rm d\,}T$ has an eigenvalue which is a root of unity. If $G$ is not nilpotent, by Lemma 3.2, ${\rm d\,}T$ has an eigenvalue which is a root of unity. We may replace $T$ by $T^{m}$, for some $m\in{\mathbb{N}}$, and assume that 1 is an eigenvalue of ${\rm d\,}T$. Let $X\in{\mathfrak{g}}$, $X\neq 0$ be such that ${\rm d\,}T(X)=X$ and let $A=\{\exp tX\}$, a one-parameter subgroup, it is either closed or its closure $\bar{A}$ is compact. Since $T|_{A}={\rm Id}$, we have that $T|_{\bar{A}}={\rm Id}$ and it follows that $\bar{A}$ normalises $C(T)$ and $\bar{A}\cap C(T)=\{e\}$.
We now show that $C(T)$ normalises $\bar{A}$, which would imply that $\bar{A}\,C(T)$ is abelian as $C(T)$ and $A$ are so. Here, $T(\bar{A})=\bar{A}$ and $\bar{A}\in{\rm Sub}^{p}_{G}$. Now for any nonzero $c\in C(T)$, we have that
$T^{n}(c\bar{A}c^{-1})=T^{n}(c)\,\bar{A}\,T^{n}(c^{-1})\to\bar{A}.$
Since $T$ acts distally on ${\rm Sub}^{p}_{G}$, we get that $c\bar{A}c^{-1}=\bar{A}$, for all $c\in C(T)$. Thus $C(T)$ normalises $\bar{A}$, and hence $\bar{A}\,C(T)$ is abelian.
Let $H=\overline{A\times C(T)}$. Then $H$ is a closed $T$-invariant abelian subgroup and $\bar{A}\subset H$. First, suppose $H$ is a vector group. By Theorem 1.1 (1), $T|_{H}\in{\mathcal{K}}{\mathcal{D}}$, for some compact subgroup ${\mathcal{K}}$ in ${\rm GL}(H)$ and the center ${\mathcal{D}}$ of ${\rm GL}(H)$, and thus all the eigenvalues of $T|_{H}$ have the same absolute value. Since $T|_{\bar{A}}={\rm Id}$ and $T|_{C(T)}$ have eigenvalues of absolute value $\alpha\neq 1$, it leads to a contradiction. Thus $C(T)$ must be trivial in this case.
Step 4: Now suppose $H$ is not a vector group. Since $C(T)$ is closed and the image of $A$ is dense in $H/C(T)$, we get that $H=K\times C(T)$ for the maximal compact subgroup $K$ of $H$. As $T$ acts trivially on $H/C(T)$ and $K$ is $T$-invariant, we get that $T|_{K}={\rm Id}$.